OpenStudy (anonymous):

if 5x + 13 = 31, what is the value of sqrt(5x + 31)?

OpenStudy (cp9454):

\[\sqrt{595}/5\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (cp9454):

find out the value of x from given equation and put in the value under sqrt

OpenStudy (anonymous):

im confused

OpenStudy (alekos):

It should be sqrt (31)

OpenStudy (anonymous):

no i dont think thats right

OpenStudy (anonymous):

but i dont know if im right

OpenStudy (anonymous):

@hartnn help me!!!!

OpenStudy (alekos):

Its the square root of 31

hartnn (hartnn):

\(\Large 5x + 13 = 31 \\ \Large \sqrt{5x + 13} =\sqrt{ 31} \)

OpenStudy (anonymous):

no its not any of the answers

hartnn (hartnn):

then please verify your question

OpenStudy (anonymous):

A.)sqrt(13) B.)sqrt(173/5) C.)7 D.)13 E.)169

OpenStudy (anonymous):

those are thee answers

hartnn (hartnn):

whats the question ?

OpenStudy (anonymous):

its wrong. i have 2 more tries

hartnn (hartnn):

please verify question, or something is wrong with the question

OpenStudy (anonymous):

i dont know what to do i know that B and A is out.

hartnn (hartnn):

you please post your question again

hartnn (hartnn):

or attach the screenshot of the question

OpenStudy (anonymous):

it papaer

hartnn (hartnn):

value of of sqrt(5x + 13) is surely sqrt 31

OpenStudy (anonymous):

summer homework

hartnn (hartnn):

either question is wrong or options are wrong

OpenStudy (anonymous):

no there isnt any choices of that

OpenStudy (anonymous):

no i asked my friends they said they got the answer but there not going to tell me. ㅜ.ㅜ

OpenStudy (anonymous):

they want me to figure it myself to exercise my brain.

OpenStudy (anonymous):

what good friends i have

hartnn (hartnn):

i think there is some problem with the question

OpenStudy (alekos):

Take photo and send it to us

OpenStudy (anonymous):

i dont know how to

hartnn (hartnn):

don't know how to take phot or don't know how to send it to us ?

OpenStudy (alekos):

Use your phone

OpenStudy (anonymous):

both

hartnn (hartnn):

you never took your photo ?

OpenStudy (anonymous):

but i can swear to openstudy that the question is legit

OpenStudy (anonymous):

no i downloaded it

OpenStudy (alekos):

Can't help you then

OpenStudy (anonymous):

oh darn

hartnn (hartnn):

if i were you i would guess 13 but thats just a guess, it could be wrong. it is very unlikely that the answer is 169

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

how did my friends get it?

hartnn (hartnn):

i cannot read minds of your friends :P

OpenStudy (anonymous):

lol that sarcasm

OpenStudy (anonymous):

ok guys thanks for your help

hartnn (hartnn):

good luck! :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have an idea but i dont think that it will work

OpenStudy (alekos):

Go option f none of the above

OpenStudy (anonymous):

i wish

OpenStudy (anonymous):

oh **** i found out!!!!

OpenStudy (anonymous):

how

OpenStudy (anonymous):

to

OpenStudy (anonymous):

do

OpenStudy (anonymous):

it!!!

hartnn (hartnn):

how ?

OpenStudy (anonymous):

ok 5x + 13 = 31 right?

OpenStudy (anonymous):

i left 5x alone 5x = 18

OpenStudy (anonymous):

and then exchanged 5x to 18

hartnn (hartnn):

you changed your question ;-; thats cheating :P

hartnn (hartnn):

sqrt(5x + 31) is obviously 7

OpenStudy (anonymous):

sorry i havent seen the mistake i was about to tell you guys but sorry.... ㅜ.ㅜ

OpenStudy (anonymous):

you guys were alot of help though

hartnn (hartnn):

hey, no problem, all good :) i am glad you got the correct answer ^_^

OpenStudy (anonymous):

thanks